OpenStudy (emanmahmoud):

A ball is thrown upward with speed 12 m/s from the top of a building. How much later must a second ball be dropped from the same starting point if it is to hit the ground at the same time as the first ball? The initial position is 24 m above the ground. thanx in advance

OpenStudy (anonymous):

:) nw we should calculate the time for first ball to reach the earth. s=ut+1/2at^2 24=-12t+(10/2)t^2 5t^2-12t-24=0 nw solve this to get the time